OpenStudy (anonymous):

graph the equation by plotting 3 points 2y=-3x+2

OpenStudy (ranga):

choose three values of x, put it in the equation and find their y values. put the points on a graph paper and using a ruler draw a straight line passing through the three points. You can choose x = 0, 1 and 2 or any other convenient values.

OpenStudy (anonymous):

2y=-3(0)+2 is this right ranga

OpenStudy (ranga):

yes, put x = 0 in 2y = -3x + 2 2y = -3(0) + 2 2y = 0 + 2 = 2 divide both sides by 2 y = 1 So one point is: (0,1) Try x = 1 and find y Then try x = 2 and find y Plot the three points and a draw a straight line through them.